OpenStudy (anonymous):

How many different ways can the students at a school select the president, vice president, and secretary from a group of 5 people?

OpenStudy (kropot72):

The order of each selection is important. Therefore the number of permutations is needed. \[5P3=\frac{5!}{(5-3)!}=you\ can\ calculate\]

OpenStudy (anonymous):

15

OpenStudy (kropot72):

Not really. \[5P3=\frac{5!}{(5-3)!}=\frac{5\times4\times3\times2\times1}{2\times1}=you\ can\ calculate\]

OpenStudy (anonymous):

60

OpenStudy (kropot72):

Correct!

OpenStudy (anonymous):

oh ok

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

thank you thank you

OpenStudy (kropot72):

yw